OpenStudy (anonymous):

2x(x-5)^4-x^2(4)(x-5)^3 -helpp! (:

OpenStudy (anonymous):

I'm not sure if you meant any of the "x"'s you put to represent multiplication ... Is this the question ? \[2*x*(x-5)^2-x^2*4*(x-5)^3\]

OpenStudy (anonymous):

Should (x-5)^4 not (x-5)^2 above ...

OpenStudy (anonymous):

no,sorry :o,the x's are variables lol

OpenStudy (anonymous):

Assuming I didn't mess up ... \[2x(x-5)^4-x^2*4*(x-5)^3\] \[=(x-5)^3*( 2x(x-5) -4x^2 )\] \[=(x-5)^3(-2x)(x+5)\] \[=-2x(x+5)(x-5)^3\]

OpenStudy (anonymous):

ahh thanks!!! :D

OpenStudy (anonymous):

yw, and click "Good Answer" :)